# Metis-SPECS: Decoupling Learning via Self-distilled Preference-based Cold Start for VLMs

### v2

@@ Abstract
Reinforcement learning (RL) with verifiable rewards has recently catalyzed a wave of “MLLM-r1” approaches that bring RL to vision language models. Most representative paradigms begin with a cold start, typically employing supervised fine-tuning (SFT), to initialize the policy before RL. However, SFT-based cold start adopts the reasoning paradigm intertwined with task solution and output format, which may induce instruction-style overfitting, weakens out-of-distribution generalization, and ultimately affects downstream RL. We revisit the cold start along two views, its training method and data construction, and introduce the Generalization Factor (GF) coefficient to quantify the generalization capability under different methods. Our empirical study finds that preference–based training methods (e.g. DPO) generalizes better than SFT-based methods in cold start. Motivated by this, we propose SPECS—a Self-distilled, Preference-based Cold Start framework that decouples multimodal learning: (1) generates introspective preference data pairs via self-distillation, avoiding reliance on larger teachers or manual annotation; (2) performs preference–based training to learn, focusing on shallow, transferable surface-form criteria (format, structure, style) rather than memorizing content; and (3) hands off to RL with verifiable rewards for deep reasoning results. Experimental results across multiple multimodal benchmarks show that our decoupling learning framework yields consistent performance gains over strong baselines, improving MEGA-Bench by 4.1% and MathVista by 12.2%. Additional experiments indicate that SPECS contributes to reducing in-distribution “stuckness,” improving exploration, stabilizing training, and raising the performance ceiling.  [^1] Project Page: https://github.com/Kwen-Chen/SPECS-VL
@@ 1 Introduction
Recently, inspired by the success of DeepSeek-R1 (guo2025deepseek), in effectively enhancing the reasoning capabilities of large models through reinforcement learning (RL) with verifiable reward (lambert2024tulu; guo2025deepseek), a growing body of work has begun to apply RL directly to vision language models (VLMs). This has led to a wave of exciting “MLLM-r1” research (meng2025mmeureka; shen2025vlm; peng2025lmm; zhou2025r1; zhang2025r1; wang2025sota; wang2025vl; zheng2025deepeyes; ma2025one), which leverage similar principles to advance multimodal reasoning.
@@
Previous research has indicated that prior to RL, employing a pre-training or warm-up phase (which is termed “cold start”), can significantly improve the readability, stability, and even the final performance of RL training (guo2025deepseek). Currently, the most commonly used cold start strategy is supervised fine-tuning (SFT) , where the model is first fine-tuned on a set of high-quality Chain-of-Thought reasoning data to provide a better initial policy for the subsequent RL phase (wei2025advancing; yang2025r1; huang2025vision; deng2025openvlthinker). This strategy enables the model to be trained on complex reasoning data during the cold start phase, thereby acquiring reasoning ability.
@@
The common understanding behind SFT-based cold start is that reasoning abilities, reasoning format and other learning objectives can be jointly learned during the cold start phase. However, such an SFT-based joint learning paradigm may largely affect the model’s generalization capability (wu2025generalization; chu2025sft), and consequently degrade subsequent RL (chen2025sft). This raises an important research issue of quantifying and improving the model’s generalization capability during cold start and working in concert with subsequent RL.
@@
To address the above limitations, we consider an alternative learning paradigm, which separates the learning process into hierarchical stages based on the idea that cold start phase focused more on shallow, surface-form learning to avoid prematurely getting stuck in in-distribution problem solving, while subsequent RL focuses on the deep-level learning of a solution to boost the overall performance (bengio2009curriculum). Thus, the intuition of our adopting decoupling learning for multimodal reasoning is that the selection of pre-training methods in cold start needs to better support the subsequent RL, both in terms of generalization and by having separate objectives to facilitate better final results.
@@
Another important issue is the generation of cold start data. Previously, the prohibitive cost of human annotation has motivated a growing body of research to explore the use of synthetic data. This often involves using a more capable large model as a “teacher” to distill data for a smaller “student” model. (zhang2025oasis; yao2024mulberry; xu2024llava; huang2025vision). However, when the capability gap between the teacher model and the student model is too large, it can lead to a decline in model performance (zhang2023towards). Alternatively, the DeepSeek-R1-Zero paradigm (guo2025deepseek) first directly applies reinforcement learning to the base model for obtaining R1-Zero and then generates cold start data by zero model itself. This paradigm has achieved very remarkable performance; yet it still has the limitation of reliance on the SFT cold start and the constraints between SFT and subsequent RL, thereby leaving room for further improvement.
@@
In this paper, to examine the suitable cold start training method, we propose the Generalization Factor (GF) coefficient in Section 2 to quantify the generalization capability of the model and conduct an empirical study to evaluate different training methods. We identify that Direct Preference Optimization (DPO) (rafailov2023direct) based on preference data is a cold start approach that enables the model to have better performance. On this basis, we present the Self-distilled Preference-based Cold-Start (SPECS) framework in Section 3. By decoupling the learning objectives during DPO to focus on output format, we create a pre-aligned model that serves as a superior starting point for the final RL fine-tuning. Our experiments show that this method leads to more stable, efficient training, and a higher performance ceiling compared to the advanced and strong baseline.
@@ 3.2 Self-Distillation for Preference Data Generation
  • Chosen Response Filtration. For the chosen response ($y_{i}^{+}$), we use Gemini-2.5 flash (comanici2025gemini) as an evaluator. Assesses whether the reasoning path in the $\pi_{GRPO-zero}$ response aligns correctly with its final answer. Only responses in which the reasoning and the answer are consistent are retained, forming a high-quality pool of candidates. For more analysis of this content, please refer to Appendix E.
@@ 3.3 DPO-based Pre-Alignment for Cold-Start
Methodology: To achieve this pre-alignment, we employ DPO (rafailov2023direct), a powerful technique that directly optimizes the language model on preference data without the need for an explicit reward model. The standard DPO loss function is defined as:
@@
To augment this process, we incorporate an SFT loss computed on the “chosen” samples, which serves as a form of regularization. It ensures that while the model learns the directional preference signal from DPO, it does not drift far from the core distribution of high-quality text embodied by the chosen responses (rao2025apt). The combined loss function is thus:
@@ 3.4 Final GRPO Fine-tuning
For the final stage of fine-tuning, we employ the GRPO algorithm (shao2024deepseekmath). This process is guided by a composite reward function that combines format and accuracy components to evaluate the model’s output, $o$, for a given question, $q$.
@@ 4.1 Experiment Settings
Dataset and Benchmark: The data utilized for training $\pi_{GRPO-zero}$ model in Stage 1 and for the final GRPO fine-tuning of the cold-started model in Stage 3 is composed of the Orsta47K (ma2025one) and virl39K (wang2025vl) datasets. In Stage 2 of cold start training, we used 9K self-distilled data. This composition is designed to enhance the model’s general and mathematical reasoning capabilities. We conduct evaluations on multiple benchmark datasets, including MEGA-Bench (chen2025mega-bench), MMMU (yue2024mmmu), MathVista (lu2024mathvista), MATH-Vision (wang2024measuring), and MathVerse (zhang2024mathverse).
@@
    The remaining results are from the MEGA-Bench Leaderboard and ma2025one.
@@
Baseline: Our comparative analysis is grounded on two primary categories of models. The first category comprises open-source general VLMs, including QwenVL-2-7B (Qwen2VL), QwenVL-2.5-7B (Qwen2.5-VL), InternVL2-8B (chen2024internvl), InternVL2.5-8B (chen2024internvl), Kimi-VL-A3B (kimiteam2025kimivltechnicalreport), and DeepSeek-VL-7B (lu2024deepseekvl). The second category focuses on models specifically engineered for advanced reasoning tasks. This group includes Kimi-VL-A3B-Thinking (kimiteam2025kimivltechnicalreport), R1-Onevision (yang2025r1), VLAA-Thinking (chen2025sft), MM-Eureka-7B (meng2025mmeureka), VL-Rethinker-7B (wang2025vl), and Orsta-7B (ma2025one).
Implementation Details: We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Qwen2.5-VL), as our base model. For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^5] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. For the DPO training in Stage 2, as well as for the comparative SFT experiments, we leverage the LlamaFactory  [^6] https://github.com/hiyouga/LLaMA-Factory framework. In this configuration, the training batch size is set to $64$, the learning rate is maintained at $1\times 10^{-6}$, and the hyperparameter $\lambda$ for the hybrid loss function is set to $1$. The prompt used during training is shown in Appendix B. Some more detailed settings can be found in Appendix F.
@@ 5 Related Work
The application of RL has emerged as a highly effective method for enhancing the reasoning capabilities of large language models, with notable successes in the text-only domain such as DeepSeek-R1 (guo2025deepseek), which leverages RLVR (lambert2024tulu; guo2025deepseek). Inspired by these advancements, a substantial and rapidly growing body of research has begun adapting RL techniques for VLMs. This has catalyzed a wave of “MLLM-r1” studies, all aiming to harness similar principles to unlock more advanced multimodal reasoning abilities. For instance, MM-Eureka (meng2025mmeureka) explores the enhancement of multimodal reasoning abilities through rule-based RL by constructing high-quality multimodal reasoning datasets. VL-Rethinker (wang2025vl) stimulates the slow thinking and self-reflection abilities of VLMs through RL. Orsta (ma2025one) establishes a unified RL system that supports VLMs in jointly learning visual reasoning and perception tasks. VLM-R1 (shen2025vlm) extends R1-style RL to VLMs for visual understanding tasks to improve their visual reasoning abilities. LMM-R1 (peng2025lmm) enhances the model’s basic reasoning ability and multimodal generalization ability through a two-stage training strategy of basic reasoning enhancement and multimodal generalization training. R1-VL (zhang2025r1) realizes the self-improvement of MLLMs’ reasoning ability by solving the sparse reward problem through Step-wise GRPO. DeepEyes (zheng2025deepeyes) motivates the model’s “Thinking with Images” ability through RL. VisualThinker-R1-Zero (zhou2025r1) performs RL directly without any supervised fine-tuning of the model to reproduce the “aha moment”.
A crucial precursor to effective RL is the “cold-start” phase, which initializes the model’s policy before the RL stage begins. The conventional strategy for this phase is SFT, a foundational step adopted by many leading models to establish a strong baseline performance (wei2025advancing; yang2025r1; huang2025vision; deng2025openvlthinker). In parallel with refining cold-start methods, the prohibitive cost of human annotation has driven the field towards synthetic data generation. This approach often involves using powerful teacher models to distill vast amounts of data for training smaller student models (zhang2025oasis; xu2024llava; huang2025vision). Vision-R1 (huang2025vision) cold-starts the model before applying RL by synthesizing 100K high-quality long CoT instructions. LLaVA-CoT (xu2024llava) integrates multiple mainstream visual question answering datasets and uses advanced large models to synthesize 99K valid image-question-answer pairs. R1-Onevision (yang2025r1) adopts a two-stage training strategy of SFT + RL by synthesizing a 155K instruction set.
@@ Appendix A The Use of Large Language Models
In preparing this manuscript, we utilized the Large Language Model (LLM) Gemini-2.5-pro (comanici2025gemini) to polishing the text. Its application was strictly limited to correcting spelling and grammatical errors. The authors manually reviewed and verified all AI-assisted modifications to ensure factual accuracy. The core ideas, methodologies, and figures presented are entirely the original work of the human authors.
@@ C.1 Preliminaries
Direct Preference Optimization. DPO(rafailov2023direct) refines models by learning directly from preference data, consisting of a prompt $x$, a chosen response $y_{c}$, and a rejected response $y_{r}$. The DPO loss function is designed to increase the relative probability of the chosen response over the rejected one: L_DPO(π_θ; π_ref) = -E_(x, y_c, y_r) ∼D [ logσ( βlogπθ(yc—x)πref(yc—x) - βlogπθ(yr—x)πref(yr—x) ) ]
@@ C.2 Analysis of Loss Balance Coefficients
Exploration Capability via Rollout Branching Factor (RBF). Beyond task performance, we investigated the intrinsic exploratory capacity of models, a critical factor for successful RL. We measure this using the RBF (deng2025trial), which quantifies the diversity of a model’s generation by counting the number of candidate tokens within the probability mass of top p (0.95) during decoding. A higher RBF signifies greater generation diversity and, consequently, a stronger capacity for exploration. As shown in Table 6, our DPO-based cold-start method yields a substantially higher RBF than the SFT baseline. This finding suggests that our approach cultivates a broader exploration space, which is highly beneficial for the subsequent RL phase, enabling the model to discover more diverse and potentially higher-quality solutions. This outcome highlights a key advantage of our method in preparing models for alignment.

### v3

@@ Abstract
Reinforcement learning (RL) with verifiable rewards has recently catalyzed a wave of “MLLM-r1” approaches that bring RL to vision language models. Most representative paradigms begin with a cold start, typically employing supervised fine-tuning (SFT), to initialize the policy before RL. However, SFT-based cold start adopts the reasoning paradigm intertwined with task solution and output format, which may induce instruction-style overfitting, weakens out-of-distribution generalization, and ultimately affects downstream RL. We revisit the cold start along two views, its training method and data construction, and introduce the Generalization Factor (GF) coefficient to quantify the generalization capability under different methods. Our empirical study finds that preference–based training methods (e.g. DPO) generalizes better than SFT-based methods in cold start. Motivated by this, we propose SPECS—a Self-distilled, Preference-based Cold Start framework that decouples multimodal learning: (1) generates introspective preference data pairs via self-distillation, avoiding reliance on larger teachers or manual annotation; (2) performs preference–based training to learn, focusing on shallow, transferable surface-form criteria (format, structure, style) rather than memorizing content; and (3) hands off to RL with verifiable rewards for deep reasoning results. Experimental results across multiple multimodal benchmarks show that our decoupling learning framework yields consistent performance gains over strong baselines, improving MEGA-Bench by 4.1% and MathVista by 12.2%. Additional experiments indicate that SPECS contributes to reducing in-distribution “stuckness,” improving exploration, stabilizing training, and raising the performance ceiling.
Project Page: https://kwen-chen.github.io/SPECS-VL/
@@ 1 Introduction
Recently, inspired by the success of DeepSeek-R1 (Guo et al., 2025), in effectively enhancing the reasoning capabilities of large models through reinforcement learning (RL) with verifiable reward (Lambert et al., 2024; Guo et al., 2025), a growing body of work has begun to apply RL directly to vision language models (VLMs). This has led to a wave of exciting “MLLM-r1” research (Meng et al., 2025; Shen et al., 2025; Peng et al., 2025; Zhou et al., 2025; Zhang et al., 2025a; Wang et al., 2025b; a; Zheng et al., 2025; Ma et al., 2025), which leverage similar principles to advance multimodal reasoning.
@@
Previous research has indicated that prior to RL, employing a pre-training or warm-up phase (which is termed “cold start”), can significantly improve the readability, stability, and even the final performance of RL training (Guo et al., 2025). Currently, the most commonly used cold start strategy is supervised fine-tuning (SFT) , where the model is first fine-tuned on a set of high-quality Chain-of-Thought reasoning data to provide a better initial policy for the subsequent RL phase (Wei et al., 2025; Yang et al., 2025; Huang et al., 2025; Deng et al., 2025b). This strategy enables the model to be trained on complex reasoning data during the cold start phase, thereby acquiring reasoning ability.
@@
The common understanding behind SFT-based cold start is that reasoning abilities, reasoning format and other learning objectives can be jointly learned during the cold start phase. However, such an SFT-based joint learning paradigm may largely affect the model’s generalization capability (Wu et al., 2025; Chu et al., 2025), and consequently degrade subsequent RL (Chen et al., 2025a). This raises an important research issue of quantifying and improving the model’s generalization capability during cold start and working in concert with subsequent RL.
@@
To address the above limitations, we consider an alternative learning paradigm, which separates the learning process into hierarchical stages based on the idea that cold start phase focused more on shallow, surface-form learning to avoid prematurely getting stuck in in-distribution problem solving, while subsequent RL focuses on the deep-level learning of a solution to boost the overall performance (Bengio et al., 2009). Thus, the intuition of our adopting decoupling learning for multimodal reasoning is that the selection of pre-training methods in cold start needs to better support the subsequent RL, both in terms of generalization and by having separate objectives to facilitate better final results.
@@
Another important issue is the generation of cold start data. Previously, the prohibitive cost of human annotation has motivated a growing body of research to explore the use of synthetic data. This often involves using a more capable large model as a “teacher” to distill data for a smaller “student” model. (Zhang et al., 2025b; Yao et al., 2024; Xu et al., 2024; Huang et al., 2025). However, when the capability gap between the teacher model and the student model is too large, it can lead to a decline in model performance (Zhang et al., 2023). Alternatively, the DeepSeek-R1-Zero paradigm (Guo et al., 2025) first directly applies reinforcement learning to the base model for obtaining R1-Zero and then generates cold start data by zero model itself. This paradigm has achieved very remarkable performance; yet it still has the limitation of reliance on the SFT cold start and the constraints between SFT and subsequent RL, thereby leaving room for further improvement.
@@
In this paper, to examine the suitable cold start training method, we propose the Generalization Factor (GF) coefficient in Section 2 to quantify the generalization capability of the model and conduct an empirical study to evaluate different training methods. We identify that Direct Preference Optimization (DPO) (Rafailov et al., 2023) based on preference data is a cold start approach that enables the model to have better performance. On this basis, we present the Self-distilled Preference-based Cold-Start (SPECS) framework in Section 3. By decoupling the learning objectives during DPO to focus on output format, we create a pre-aligned model that serves as a superior starting point for the final RL fine-tuning. Our experiments show that this method leads to more stable, efficient training, and a higher performance ceiling compared to the advanced and strong baseline.
@@ 3.2 Self-Distillation for Preference Data Generation
  • Chosen Response Filtration. For the chosen response ($y_{i}^{+}$), we use Gemini-2.5 flash (Comanici et al., 2025) as an evaluator. Assesses whether the reasoning path in the $\pi_{GRPO-zero}$ response aligns correctly with its final answer. Only responses in which the reasoning and the answer are consistent are retained, forming a high-quality pool of candidates. For more analysis of this content, please refer to Appendix E.
@@ 3.3 DPO-based Pre-Alignment for Cold-Start
Methodology: To achieve this pre-alignment, we employ DPO (Rafailov et al., 2023), a powerful technique that directly optimizes the language model on preference data without the need for an explicit reward model. The standard DPO loss function is defined as:
@@
To augment this process, we incorporate an SFT loss computed on the “chosen” samples, which serves as a form of regularization. It ensures that while the model learns the directional preference signal from DPO, it does not drift far from the core distribution of high-quality text embodied by the chosen responses (Rao et al., 2025). The combined loss function is thus:
@@ 3.4 Final GRPO Fine-tuning
For the final stage of fine-tuning, we employ the GRPO algorithm (Shao et al., 2024). This process is guided by a composite reward function that combines format and accuracy components to evaluate the model’s output, $o$, for a given question, $q$.
@@ 4.1 Experiment Settings
Dataset and Benchmark: The data utilized for training $\pi_{GRPO-zero}$ model in Stage 1 and for the final GRPO fine-tuning of the cold-started model in Stage 3 is composed of the Orsta47K (Ma et al., 2025) and virl39K (Wang et al., 2025a) datasets. In Stage 2 of cold start training, we used 9K self-distilled data. This composition is designed to enhance the model’s general and mathematical reasoning capabilities. We conduct evaluations on multiple benchmark datasets, including MEGA-Bench (Chen et al., 2025b), MMMU (Yue et al., 2024), MathVista (Lu et al., 2024b), MATH-Vision (Wang et al., 2024a), and MathVerse (Zhang et al., 2024).
@@
    The remaining results are from the MEGA-Bench Leaderboard and Ma et al. (2025).
@@
Baseline: Our comparative analysis is grounded on two primary categories of models. The first category comprises open-source general VLMs, including QwenVL-2-7B (Wang et al., 2024b), QwenVL-2.5-7B (Bai et al., 2025), InternVL2-8B (Chen et al., 2024), InternVL2.5-8B (Chen et al., 2024), Kimi-VL-A3B (Team et al., 2025), and DeepSeek-VL-7B (Lu et al., 2024a). The second category focuses on models specifically engineered for advanced reasoning tasks. This group includes Kimi-VL-A3B-Thinking (Team et al., 2025), R1-Onevision (Yang et al., 2025), VLAA-Thinking (Chen et al., 2025a), MM-Eureka-7B (Meng et al., 2025), VL-Rethinker-7B (Wang et al., 2025a), and Orsta-7B (Ma et al., 2025).
Implementation Details: We utilize the open-source Multimodal Large Language Model, Qwen2.5-VL-7B (Bai et al., 2025), as our base model. For the GRPO training in Stage 1 and Stage 3, we employ the MM-EUREKA  [^4] https://github.com/ModalMinds/MM-EUREKA framework. The rollout and training batch sizes are both set to $128$, with $8$ rollouts generated per sample. The learning rate is configured to $1\times 10^{-6}$. For the DPO training in Stage 2, as well as for the comparative SFT experiments, we leverage the LlamaFactory  [^5] https://github.com/hiyouga/LLaMA-Factory framework. In this configuration, the training batch size is set to $64$, the learning rate is maintained at $1\times 10^{-6}$, and the hyperparameter $\lambda$ for the hybrid loss function is set to $1$. The prompt used during training is shown in Appendix B. Some more detailed settings can be found in Appendix F.
@@ 5 Related Work
The application of RL has emerged as a highly effective method for enhancing the reasoning capabilities of large language models, with notable successes in the text-only domain such as DeepSeek-R1 (Guo et al., 2025), which leverages RLVR (Lambert et al., 2024; Guo et al., 2025). Inspired by these advancements, a substantial and rapidly growing body of research has begun adapting RL techniques for VLMs. This has catalyzed a wave of “MLLM-r1” studies, all aiming to harness similar principles to unlock more advanced multimodal reasoning abilities. For instance, MM-Eureka (Meng et al., 2025) explores the enhancement of multimodal reasoning abilities through rule-based RL by constructing high-quality multimodal reasoning datasets. VL-Rethinker (Wang et al., 2025a) stimulates the slow thinking and self-reflection abilities of VLMs through RL. Orsta (Ma et al., 2025) establishes a unified RL system that supports VLMs in jointly learning visual reasoning and perception tasks. VLM-R1 (Shen et al., 2025) extends R1-style RL to VLMs for visual understanding tasks to improve their visual reasoning abilities. LMM-R1 (Peng et al., 2025) enhances the model’s basic reasoning ability and multimodal generalization ability through a two-stage training strategy of basic reasoning enhancement and multimodal generalization training. R1-VL (Zhang et al., 2025a) realizes the self-improvement of MLLMs’ reasoning ability by solving the sparse reward problem through Step-wise GRPO. DeepEyes (Zheng et al., 2025) motivates the model’s “Thinking with Images” ability through RL. VisualThinker-R1-Zero (Zhou et al., 2025) performs RL directly without any supervised fine-tuning of the model to reproduce the “aha moment”.
A crucial precursor to effective RL is the “cold-start” phase, which initializes the model’s policy before the RL stage begins. The conventional strategy for this phase is SFT, a foundational step adopted by many leading models to establish a strong baseline performance (Wei et al., 2025; Yang et al., 2025; Huang et al., 2025; Deng et al., 2025b). In parallel with refining cold-start methods, the prohibitive cost of human annotation has driven the field towards synthetic data generation. This approach often involves using powerful teacher models to distill vast amounts of data for training smaller student models (Zhang et al., 2025b; Xu et al., 2024; Huang et al., 2025). Vision-R1 (Huang et al., 2025) cold-starts the model before applying RL by synthesizing 100K high-quality long CoT instructions. LLaVA-CoT (Xu et al., 2024) integrates multiple mainstream visual question answering datasets and uses advanced large models to synthesize 99K valid image-question-answer pairs. R1-Onevision (Yang et al., 2025) adopts a two-stage training strategy of SFT + RL by synthesizing a 155K instruction set.
@@ Appendix A The Use of Large Language Models
In preparing this manuscript, we utilized the Large Language Model (LLM) Gemini-2.5-pro (Comanici et al., 2025) to polishing the text. Its application was strictly limited to correcting spelling and grammatical errors. The authors manually reviewed and verified all AI-assisted modifications to ensure factual accuracy. The core ideas, methodologies, and figures presented are entirely the original work of the human authors.
@@ C.1 Preliminaries
Direct Preference Optimization. DPO(Rafailov et al., 2023) refines models by learning directly from preference data, consisting of a prompt $x$, a chosen response $y_{c}$, and a rejected response $y_{r}$. The DPO loss function is designed to increase the relative probability of the chosen response over the rejected one: L_DPO(π_θ; π_ref) = -E_(x, y_c, y_r) ∼D [ logσ( βlogπθ(yc—x)πref(yc—x) - βlogπθ(yr—x)πref(yr—x) ) ]
@@ C.2 Analysis of Loss Balance Coefficients
Exploration Capability via Rollout Branching Factor (RBF). Beyond task performance, we investigated the intrinsic exploratory capacity of models, a critical factor for successful RL. We measure this using the RBF (Deng et al., 2025a), which quantifies the diversity of a model’s generation by counting the number of candidate tokens within the probability mass of top p (0.95) during decoding. A higher RBF signifies greater generation diversity and, consequently, a stronger capacity for exploration. As shown in Table 6, our DPO-based cold-start method yields a substantially higher RBF than the SFT baseline. This finding suggests that our approach cultivates a broader exploration space, which is highly beneficial for the subsequent RL phase, enabling the model to discover more diverse and potentially higher-quality solutions. This outcome highlights a key advantage of our method in preparing models for alignment.